# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 2.3. Irregular types
Let us spell out the definition of irregular type. Recall that $\mathscr{T}_{\Sigma,\sigma}^{\leq k}=\mathcal{O}_{\Sigma}(k\sigma(B))/\mathcal{O}_{\Sigma}$. If $(e_{1},\ldots,e_{r})$ is a basis of $\mathfrak{t}$ as a complex vector space, then every global section $Q$ of $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ can be written uniquely as $\sum_{j=1}^{r}Q^{(j)}e_{j}$, with $Q^{(j)}$ a section of $\mathcal{O}_{\Sigma}(k\sigma(B))/\mathcal{O}_{\Sigma}$. Such a section is a meromorphic function defined locally (around each point of $\sigma(B)$, cf. section 2.2.2) up to holomorphic terms, and with pole along $\sigma(B)$ of order at most $k$. Therefore, irregular types can be thought of (locally) as meromorphic $\mathfrak{t}$-valued functions with pole along $\sigma(B)$ of order at most $k$, up to holomorphic terms.
@@
$\mathscr{T}_{\Sigma,\sigma}^{\leq k}\rightarrow\psi_{*}\psi^{*}\mathscr{T}_{\Sigma,\sigma}^{\leq k}\rightarrow\psi_{*}\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}$
@@
Let $\pi\colon\Sigma\rightarrow B$ be a Riemann surface over $B$ of genus $g$, $\sigma\colon B\rightarrow\Sigma$ a section and $k\geq 1$ an integer. The functor $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ is representable by a vector bundle of rank $rk$ on $B$.
@@
where the last identification follows from Lemma 2.2.8. Hence, the functor $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ is the functor of global sections of $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{|B}^{\leq k}$, which is a locally free sheaf of rank $rk$ over $B$ by Lemma 2.2.4. The vector bundle over $B$ attached to this locally free sheaf represents $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ (see section 4.1.1). ∎
2.3.3.
We call the vector bundle in the previous proposition the bundle of irregular types on $B$ with pole at $\sigma$ of order at most $k$, and we still denote it by $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$. We also denote by $\mathscr{T}^{\leq k}_{B,\sigma}$ the vector bundle of rank $k$ over $B$ representing the functor sending $B^{\prime}\rightarrow B$ to $\Gamma(\Sigma^{\prime},\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k})$. If $(\sigma_{1},\ldots,\sigma_{m})$ are non-intersecting sections, we denote by $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma_{1},\ldots,\sigma_{m}}$ the fibre product of the bundles $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma_{i}}$ over $B$.
@@ 2.3.4. The root stratification
We will stratify the space $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ using the following inputs.
@@
    For $j\leq k$ we have natural injections $\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq j}\rightarrow\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}$ for every $B^{\prime}\rightarrow B$ apply the snake lemma to
    00 0 $\rightarrow\mathcal{O}_{\Sigma^{\prime}}\rightarrow\mathcal{O}_{\Sigma^{\prime}}(j\sigma(B^{\prime}))\rightarrow\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq j}\rightarrow 0$
    00 0 $\rightarrow\mathcal{O}_{\Sigma^{\prime}}\rightarrow\mathcal{O}_{\Sigma^{\prime}}(k\sigma(B^{\prime}))\rightarrow\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}\rightarrow 0.$
    Hence we obtain a morphism $\mathscr{T}^{\leq j}_{B,\sigma}\rightarrow\mathscr{T}^{\leq k}_{B,\sigma}$ identifying the source with a sub-vector bundle of the target (similarly, we have a morphism $\mathscr{I\!\!T}^{\mathfrak{t},\leq j}_{B,\sigma}\rightarrow\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$).
@@
    A root $\alpha\in\Phi$ induces natural surjections of sheaves $\mathfrak{t}\otimes\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}\rightarrow\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}$ for every $B^{\prime}\rightarrow B$, hence a submersion $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}\rightarrow\mathscr{T}^{\leq k}_{B,\sigma}$ of vector bundles over $B$.
Given a collection of integers $\mathbf{d}=(d_{\alpha})_{\alpha\in\Phi}$ with $d_{\alpha}\leq k$ for every $\alpha\in\Phi$, we denote by $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\leq\mathbf{d}}_{B,\sigma}$ the intersection of the fibre products $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}\times_{\mathscr{T}^{\leq k}_{B,\sigma}}\mathscr{T}^{\leq d_{\alpha}}_{B,\sigma}\subset\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ as $\alpha\in\Phi$ varies. This is a closed submanifold of $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ (we will justify in section 2.3.5 why it is a manifold); by construction, it is the moduli space of irregular types at $\sigma$ with pole order bounded by $k$, and such that after evaluation at each $\alpha\in\Phi$ the pole order is bounded by $d_{\alpha}$.
@@
By construction, $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\mathbf{d}}_{B,\sigma}$ is an open submanifold of $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\leq\mathbf{d}}$, and a locally closed submanifold of $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$. It parametrises irregular types at $\sigma$ with pole order bounded by $k$, and such that after evaluation at $\alpha\in\Phi$ the pole order is (everywhere) equal to $d_{\alpha}$.
@@ 2.3.5. Local structure of root strata
Let us describe $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\mathbf{d}}_{B,\sigma}$ locally on $B$. Given $b\in B$, we choose small enough open polydiscs $U\ni b$ and $V\ni\sigma(b)$ with charts as in section 2.1.2. Restricted to $U$, the sheaf $\mathscr{T}^{\leq k}_{|B}$ is free over $\mathcal{O}_{U}$, with basis $(z_{n+1}^{-1},\ldots,z_{n+1}^{-k})$; we may, and will, use this basis to identify $\mathscr{T}^{\leq k}_{|B}$ restricted to $U$ with $\mathcal{O}_{U}^{k}$. Concretely, the point is that a meromorphic function on $V$ with pole of order at most $k$ along $\sigma(B)$ can be written uniquely, up to holomorphic terms, as a sum $f_{1}z_{n+1}^{-1}+\ldots+f_{k}z_{n+1}^{-k}$, where $f_{1},\ldots,f_{k}$ are holomorphic functions of $z_{1},\ldots,z_{n}$, which are identified with functions on $U$.
@@ 2.4. Stacks of wild Riemann surfaces
In this section we will define stacks of wild Riemann surfaces (of fixed genus, number of marked points, and maximal pole order of the irregular types). We start by introducing (families of) wild Riemann surfaces. Later on we will ask, as in [Boa14a, Definition 10.1], such families to be admissible; this will lead to a stratification of the stack of wild Riemann surfaces, mirroring the stratification on $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ described above.
@@
A wild Riemann surface over a complex manifold $B$ of genus $g$, with $m$ marked points and pole orders bounded by $p$, is a triple $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ consisting of
@@
    an $m$-tuple $\underline{a}=(a_{1},\ldots,a_{m})$ of mutually non-intersecting sections $a_{i}\colon B\rightarrow\Sigma$ of $\pi$;
@@
    an $m$-tuple $\underline{Q}=(Q_{1},\ldots,Q_{m})$ of irregular types $Q_{i}\in\Gamma(\Sigma,\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{\Sigma,a_{i}}^{\leq p})$.
2.4.2. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$
@@
For every complex manifold $B$, the groupoid $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)$ is by definition the category whose objects are wild Riemann surfaces over $B$ of genus $g$, with $m$ marked points and pole orders bounded by $p$. An isomorphism from $(\pi_{1}\colon\Sigma_{1}\rightarrow B,\underline{a}_{1},\underline{Q}_{1})$ to $(\pi_{2}\colon\Sigma_{2}\rightarrow B,\underline{a}_{2},\underline{Q}_{2})$ is an isomorphism $\psi:\Sigma_{1}\rightarrow\Sigma_{2}$ such that $\pi_{2}\circ\psi=\pi_{1}$, the pullback of $\underline{a}_{2}$ is $\underline{a}_{1}$ (in other words, $\psi\circ a_{1,i}=a_{2,i}$ for $1\leq i\leq m$) and the pullback of $\underline{Q}_{2}$ is $\underline{Q}_{1}$. As above we see that any holomorphic map $B^{\prime}\rightarrow B$ induces a functor from $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)$ to $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B^{\prime})$. Forgetting irregular types we get functorial maps $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)\rightarrow\mathsf{M}_{g,m}(B)$ for every manifold $B$, where $\mathsf{M}_{g,m}(B)$ is the groupoid of genus $g$ Riemann surfaces over $B$ with $m$ mutually disjoint sections.
2.4.3. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$
@@
Let us now define admissible families wild Riemann surfaces over a manifold $B$, following [Boa14a, Definition 10.1]. Fix integers $0\leq d_{\alpha,i}\leq p$ for each $\alpha\in\Phi$ and $1\leq i\leq m$, and let $\underline{\mathbf{d}}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{m})$ with $\mathbf{d}_{i}=(d_{\alpha,i})_{\alpha\in\Phi}$.
@@
If $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ is a wild Riemann surface as above, then for every $\alpha\in\Phi$ and every $1\leq i\leq m$ we can look at $\alpha\circ Q_{i}\in\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq p})$. Recall that we have injections of sheaves $\mathscr{T}_{\Sigma,a_{i}}^{\leq d_{\alpha,i}}\rightarrow\mathscr{T}_{\Sigma,a_{i}}^{\leq p}$ (cf. section 2.3.4). We will use the following terminology.
@@
  • We say that $\alpha\circ Q_{i}$ has order at most $d_{\alpha,i}$ if it belongs to $\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq d_{\alpha,i}})$.
@@
  • We say that $\alpha\circ Q_{i}$ has order $d_{\alpha,i}$ if it belongs to $\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq d_{\alpha,i}})$ and, for every $b\in B$, its pullback to the fibre $\Sigma_{b}=\pi^{-1}(b)$ does not belong to $\Gamma(\Sigma_{b},\mathscr{T}_{\Sigma_{b},a_{i,b}}^{\leq d_{\alpha,i}-1})$.
@@
  • We say that $Q_{i}$ has root order at most $\mathbf{d}_{i}$ (resp. root order $\mathbf{d}_{i}$) if $\alpha\circ Q_{i}$ has order at most (resp. equal to) $d_{\alpha,i}$ for every $\alpha\in\Phi$. We say that $\underline{Q}$ has root order at most $\underline{\mathbf{d}}$ (resp. root order $\underline{\mathbf{d}}$) if $Q_{i}$ has root order at most $\mathbf{d}_{i}$ (resp. order $\mathbf{d}_{i}$) for all $1\leq i\leq m$.
@@
  • We say that $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ is admissible of root order $\underline{\mathbf{d}}$ if $\underline{Q}$ has root order $\underline{\mathbf{d}}$. The groupoid of admissible wild Riemann surfaces (over $B$, of genus $g$, with $m$ marked points and pole order bounded by $p$) of root order $\underline{\mathbf{d}}$ is denoted by $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}(B)$.
@@
Let us rephrase more concretely the condition that $\alpha\circ Q_{i}$ has order $d_{\alpha,i}$. For every $b\in B$, pulling back $\alpha\circ Q_{i}\in\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq p})$ to the fibre $\Sigma_{b}$ we obtain an element $(\alpha\circ Q_{i})_{b}\in\Gamma(\Sigma_{b},\mathscr{T}_{\Sigma_{b},a_{i(b)}}^{\leq p})\simeq z^{-p}\mathcal{O}_{\Sigma_{b},a_{i(b)}}/\mathcal{O}_{\Sigma_{b},a_{i(b)}}$, where $z$ is a local coordinate on $\Sigma_{b}$ around $a_{i(b)}$. Then $\alpha\circ Q_{i}$ has order $d_{\alpha,i}$ if and only if for every $b\in B$ the pole order of $(\alpha\circ Q_{i})_{b}$ is $d_{\alpha,i}$.
@@
The assignment $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\colon\mathrm{Man}_{\mathbf{C}}\rightarrow\mathrm{Groupoids}$ sending $B$ to $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)$ is a stack.

### v2

@@ 2.3.1. Irregular type of a connection on a framed bundle
Let us spell out the above definition of irregular type for $k\geq 1$.
Recall that $\mathscr{T}_{\Sigma,\sigma}^{\leq k}=\mathcal{O}_{\Sigma}(k\sigma)/\mathcal{O}_{\Sigma}$. If $(e_{1},\ldots,e_{r})$ is a basis of $\mathfrak{t}$ as a complex vector space, then every global section $Q$ of $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ can be written uniquely as $\sum_{j=1}^{r}e_{j}\otimes Q^{(j)}$, with $Q^{(j)}$ a section of $\mathcal{O}_{\Sigma}(k\sigma)/\mathcal{O}_{\Sigma}$. Such a section is a meromorphic function defined locally (around each point of $\sigma(B)$, cf. 2.2.2) up to holomorphic terms, and with pole along $\sigma(B)$ of order at most $k$. Therefore, irregular types can be thought of (locally) as meromorphic $\mathfrak{t}$-valued functions with pole along $\sigma(B)$ of order at most $k$, up to holomorphic terms.
2.3.3.
@@
$\mathscr{T}_{\Sigma,\sigma}^{\leq k}\rightarrow\psi_{*}\psi^{*}\mathscr{T}_{\Sigma,\sigma}^{\leq k}\rightarrow\psi_{*}\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k},$
@@
Let $\pi\colon\Sigma\rightarrow B$ be a Riemann surface over $B$ of genus $g$, $\sigma\colon B\rightarrow\Sigma$ a section and $k\geq 0$ an integer. The functor $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ is representable by a vector bundle (of rank $rk$) on $B$.
@@
where the last identification follows from Lemma 2.2.8. Hence, the functor $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ is the functor of global sections of $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{|B}^{\leq k}$, which is a locally free sheaf of rank $rk$ over $B$ by Lemma 2.2.4. The vector bundle over $B$ attached to this locally free sheaf represents $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ (see section A.1.1). ∎
2.3.5.
We call the vector bundle of Proposition 2.3.4 the bundle of irregular types on $B$ with pole at $\sigma$ of order at most $k$, and we still denote it by $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$. We denote by $\mathscr{T}^{\leq k}_{B,\sigma}$ the vector bundle of rank $k$ over $B$ representing the functor sending $B^{\prime}\rightarrow B$ to $\Gamma(\Sigma^{\prime},\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k})$.
If $\underline{\sigma}=(\sigma_{1},\ldots,\sigma_{m})$ is an $m$-tuple of non-intersecting sections of $\Sigma\rightarrow B$ and $\underline{k}=(k_{1},\ldots,k_{m})\in\mathbf{Z}_{\geq 0}^{m}$, we denote by $\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{k}}_{B,\underline{\sigma}}$ the fibre product of the bundles $\mathscr{I\!\!T}^{\mathfrak{t},\leq k_{i}}_{B,\sigma_{i}}$ over $B$, for $1\leq i\leq m$.
@@ 2.3.6. The root stratification
Given $\pi\colon\Sigma\rightarrow B$ with a section $\sigma$, and an integer $k\geq 0$, we will stratify the space $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ using the following inputs:
@@
    for $0\leq j\leq k$ we have natural injections $\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq j}\rightarrow\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}$ for every $B^{\prime}\rightarrow B$, hence we obtain a morphism $\mathscr{T}^{\leq j}_{B,\sigma}\rightarrow\mathscr{T}^{\leq k}_{B,\sigma}$ identifying the source with a vector sub-bundle of the target;
@@
    similarly, we have a morphism $\mathscr{I\!\!T}^{\mathfrak{t},\leq j}_{B,\sigma}\rightarrow\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$;
  3. (3)
    a root $\alpha\in\Phi$ induces natural surjections of sheaves $\mathfrak{t}\otimes\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}\rightarrow\mathscr{T}_{\Sigma^{\prime},\sigma^{\prime}}^{\leq k}$ for every $B^{\prime}\rightarrow B$; this yields a submersion $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}\rightarrow\mathscr{T}^{\leq k}_{B,\sigma}$ of vector bundles over $B$.
Given a collection of integers $\mathbf{d}=(d_{\alpha})_{\alpha\in\Phi}$ with $0\leq d_{\alpha}\leq k$ for every $\alpha\in\Phi$, we denote by $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\leq\mathbf{d}}_{B,\sigma}$ the intersection of the fibre products $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}\times_{\mathscr{T}^{\leq k}_{B,\sigma}}\mathscr{T}^{\leq d_{\alpha}}_{B,\sigma}\subset\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ as $\alpha\in\Phi$ varies. This is a closed submanifold of $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ (we will justify in section 2.3.7 that it is a manifold); by construction, it is the moduli space of irregular types with pole at $\sigma$ of order bounded by $k$, and such that after evaluation at each $\alpha\in\Phi$ the pole order is bounded by $d_{\alpha}$.
@@
By construction, $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\mathbf{d}}_{B,\sigma}$ is an open submanifold of $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\leq\mathbf{d}}_{B,\sigma}$, and a locally closed submanifold of $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$. It parametrises irregular types with pole at $\sigma$ of order at most $k$, and such that after evaluation at $\alpha\in\Phi$ the pole order is (everywhere) equal to $d_{\alpha}$.
@@ 2.3.7. Local structure of root strata
Let us describe $\mathscr{I\!\!T}^{\mathfrak{t},\leq k,\mathbf{d}}_{B,\sigma}\rightarrow B$ locally on $B$.
Given $b\in B$, we choose small enough opens $U\ni b$ and $V\ni\sigma(b)$ with charts as in section 2.1.2. Restricted to $U$, the sheaf $\mathscr{T}^{\leq k}_{|B}$ is free over $\mathcal{O}_{U}$, with basis $(z_{n+1}^{-1},\ldots,z_{n+1}^{-k})$; let us use this basis to identify $\mathscr{T}^{\leq k}_{|B}$ restricted to $U$ with $\mathcal{O}_{U}^{k}$. Concretely, the point is that a meromorphic function on $V$ with pole of order at most $k$ along $\sigma(B)$ can be written uniquely, up to holomorphic terms, as a sum $f_{1}z_{n+1}^{-1}+\ldots+f_{k}z_{n+1}^{-k}$, where $f_{1},\ldots,f_{k}$ are holomorphic functions of the variables $z_{1},\ldots,z_{n}$, which are identified with functions on $U$.
@@ 2.4. Stacks of wild Riemann surfaces
In this section we will define stacks of (labelled, untwisted) wild Riemann surfaces, of fixed genus, number of marked points, and maximal pole orders of the irregular types. We start by introducing (families of) wild Riemann surfaces. Later on we will ask, as in [Boa14a, Definition 10.1], such families to be admissible; this will lead to a stratification of the stack of wild Riemann surfaces, mirroring the stratification on $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{B,\sigma}$ described above.
Recall that we fixed an integer $m\geq 1$ and an $m$-tuple $\underline{p}=(p_{1},\ldots,p_{m})\in\mathbf{Z}_{\geq 0}^{m}$.
@@
A wild Riemann surface over a complex manifold $B$ of genus $g$, with $m$ marked points and pole orders bounded by $\underline{p}$, is a triple $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ consisting of:
@@
    an $m$-tuple $\underline{a}=(a_{1},\ldots,a_{m})$ of mutually non-intersecting holomorphic sections $a_{i}\colon B\rightarrow\Sigma$ of $\pi$;
@@
    an $m$-tuple $\underline{Q}=(Q_{1},\ldots,Q_{m})$ of irregular types $Q_{i}\in\Gamma(\Sigma,\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{\Sigma,a_{i}}^{\leq p_{i}})$.
In particular, note that we have $Q_{i}=0$ whenever $p_{i}=0$.
2.4.2. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$
@@
For every complex manifold $B$, the groupoid $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)$ is the category whose objects are wild Riemann surfaces over $B$ of genus $g$, with $m$ marked points and pole orders bounded by $\underline{p}$. An isomorphism from $(\pi_{1}\colon\Sigma_{1}\rightarrow B,\underline{a}_{1},\underline{Q}_{1})$ to $(\pi_{2}\colon\Sigma_{2}\rightarrow B,\underline{a}_{2},\underline{Q}_{2})$ is an isomorphism $\psi\colon\Sigma_{1}\rightarrow\Sigma_{2}$ such that $\pi_{2}\circ\psi=\pi_{1}$, the pullback of $\underline{a}_{2}$ is $\underline{a}_{1}$ (in other words, $\psi\circ a_{1,i}=a_{2,i}$ for $1\leq i\leq m$) and the pullback of $\underline{Q}_{2}$ is $\underline{Q}_{1}$.
As above we see that any holomorphic map $B^{\prime}\rightarrow B$ induces a functor from $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)$ to $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B^{\prime})$. Moreover, forgetting irregular types we get functorial maps $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)\rightarrow\mathsf{M}_{g,m}(B)$ for every manifold $B$, where $\mathsf{M}_{g,m}(B)$ is the groupoid of genus $g$ Riemann surfaces over $B$ with $m$ mutually disjoint sections.
2.4.3. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$
@@
Let us now define admissible families of wild Riemann surfaces over a manifold $B$. Fix integers $0\leq d_{\alpha,i}\leq p_{i}$ for each $\alpha\in\Phi$ and $1\leq i\leq m$, and let $\underline{\mathbf{d}}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{m})$, where $\mathbf{d}_{i}=(d_{\alpha,i})_{\alpha\in\Phi}$.
@@
If $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ is a wild Riemann surface as above, then for every $\alpha\in\Phi$ and $1\leq i\leq m$ we can look at $\alpha\circ Q_{i}\in\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq p_{i}})$ - usually called an exponential factor. Recall that we have injections of sheaves $\mathscr{T}_{\Sigma,a_{i}}^{\leq d_{\alpha,i}}\rightarrow\mathscr{T}_{\Sigma,a_{i}}^{\leq p_{i}}$ (cf. section 2.3.6). We will use the following terminology:
@@
  • we say that $\alpha\circ Q_{i}$ has order at most $d_{\alpha,i}$ if it belongs to $\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq d_{\alpha,i}})$;
@@
  • we say that $\alpha\circ Q_{i}$ has order $d_{\alpha,i}$ if it belongs to $\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq d_{\alpha,i}})$ and, for every $b\in B$, its pullback to the fibre $\Sigma_{b}=\pi^{-1}(b)$ does not belong to $\Gamma(\Sigma_{b},\mathscr{T}_{\Sigma_{b},a_{i}(b)}^{\leq d_{\alpha,i}-1})$;
@@
  • we say that $Q_{i}$ has root order at most $\mathbf{d}_{i}$ (resp. root order $\mathbf{d}_{i}$) if $\alpha\circ Q_{i}$ has order at most (resp. equal to) $d_{\alpha,i}$ for every $\alpha\in\Phi$;
  • we say that $\underline{Q}$ has root order at most $\underline{\mathbf{d}}$ (resp. root order $\underline{\mathbf{d}}$) if $Q_{i}$ has root order at most $\mathbf{d}_{i}$ (resp. root order $\mathbf{d}_{i}$) for all $1\leq i\leq m$;
  • we say that $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ is admissible of root order $\underline{\mathbf{d}}$ if $\underline{Q}$ has root order $\underline{\mathbf{d}}$.
@@
The groupoid of admissible wild Riemann surfaces (over $B$, of genus $g$, with $m$ marked points and pole orders bounded by $\underline{p}$) of root order $\underline{\mathbf{d}}$ is denoted by $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}(B)$.
@@
Let us rephrase more concretely the condition that $\alpha\circ Q_{i}$ has order $d_{\alpha,i}$.
For every $b\in B$, pulling back $\alpha\circ Q_{i}\in\Gamma(\Sigma,\mathscr{T}_{\Sigma,a_{i}}^{\leq p_{i}})$ to the fibre $\Sigma_{b}$ we obtain an element $(\alpha\circ Q_{i})_{b}\in\Gamma(\Sigma_{b},\mathscr{T}_{\Sigma_{b},a_{i}(b)}^{\leq p_{i}})\simeq z^{-p_{i}}\mathcal{O}_{\Sigma_{b},a_{i}(b)}/\mathcal{O}_{\Sigma_{b},a_{i}(b)}$, where $z$ is a local coordinate on $\Sigma_{b}$ around $a_{i}(b)$. Then $\alpha\circ Q_{i}$ has order $d_{\alpha,i}$ if and only if for every $b\in B$ the pole order of $(\alpha\circ Q_{i})_{b}$ is $d_{\alpha,i}$.
@@
The assignment $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\colon\mathrm{Man}_{\mathbf{C}}\rightarrow\mathrm{Groupoids}$ sending $B$ to $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)$ is a stack.